# Investigating the Impact of Trust in Multi-Human Multi-Robot Task Allocation

### v1

@@ Abstract
Trust is essential in human-robot collaboration. Even more so in multi-human multi-robot teams where trust is vital to ensure teaming cohesion in complex operational environments. Yet, at the moment, trust is rarely considered a factor during task allocation and reallocation in algorithms used in multi-human, multi-robot collaboration contexts. Prior work on trust in single-human-robot interaction has identified that including trust as a parameter in human-robot interaction significantly improves both performance outcomes and human experience with robotic systems. However, very little research has explored the impact of trust in multi-human multi-robot collaboration, specifically in the context of task allocation. In this paper, we introduce a new trust model, the Expectation Comparison Trust (ECT) model, and employ it with three trust models from prior work and a baseline no-trust model to investigate the impact of trust on task allocation outcomes in multi-human multi-robot collaboration. Our experiment involved different team configurations, including 2 humans, 2 robots, 5 humans, 5 robots, and 10 humans, 10 robots. Results showed that using trust-based models generally led to better task allocation outcomes in larger teams (10 humans and 10 robots) than in smaller teams. We discuss the implications of our findings and provide recommendations for future work on integrating trust as a variable for task allocation in multi-human, multi-robot collaboration.
@@ I Introduction
Trust is crucial in several human-robot collaboration (HRC) tasks, such as search and rescue operations [1, 2], disaster response [3, 4], industrial collaboration [5, 6, 7, 8], and other contexts requiring multi-agent robot collaboration [9, 10, 11, 12, 13, 14, 15]. To capture the different dimensions of trust in HRC, researchers have provided different definitions of trust, including characterizing it as the belief that a human or robot team member will perform as expected in a given task scenario [16]. The guarantee that an agent possesses enough capacity to complete a task in uncertain and dynamic environments [17]. The expectation that agents collaborating on a task or project will not harm each other [18]. Other researchers have also characterized it as the belief that collaborating agents will be willing to act on the information and recommendations given to them to enable them to complete a task [16]. In this research, we align with the definition that posits that trust is the expectation that a team member has the required and expected capacity to complete a task.
@@
Prior work by researchers has adopted different computational strategies for modeling trust, including using probabilistic, non-probabilistic, and graph-based approaches to capture and model the different dynamic and behavioral aspects of trust. However, although several research studies have explored trust in single-HRC, very limited work has explored trust in the context of multi-human multi-robot (MH-MR) collaboration [19, 20, 21, 22]. Considering the increasing proliferation of larger, more diverse MH-MR teams and collaborative activities, it is crucial to investigate how different trust models perform in complex, multi-agent task allocation scenarios and their value in enhancing task allocation outcomes in MH-MR collaboration that involve complex and sensitive tasks.
In this research, we investigated the impact of trust on MH-MR task allocation performance outcomes. Our investigation involved five trust models, including a no-trust baseline model, three models from prior work, Monir et al.[23], the Xu and Dudek Model[24], the Guo and Yang Model [25], and the Expectation Confirmation Trust (ECT) Model we introduce in this research. Our ECT Model draws inspiration from the expectation confirmation theory [26] and applies its core concept for modeling trust in MH-MR contexts. Specifically, the ECT model, as illustrated in Fig. 1, posits that trust is the interplay between the level of divergence in prior expectations of performance and actual performance results, focusing especially on human expectations of robot performance. By conducting this investigation, we seek to highlight the importance of trust in MH-MR contexts and provide insights into the strengths and limitations of different trust modeling approaches in MH-MR contexts. Our research question included: How do different trust models affect task allocation in teams of varying size and composition like 2 humans and 2 robots (2H-2R), 5 humans and 5 robots (5H-5R), and 10 humans and 10 robots (10H-10R)?
Through our research, we make the following contributions: 1) We propose the ECT model to provide robotics researchers and practitioners with the ability to model trust in MH-MR task allocation; 2) We conduct an experiment to investigate the impact of introducing trust in MH-MR teams and found that trust improved the performance of higher team configurations; and 3) We further show that there is no monolithic trust model, as other models performed well in the experiment. This finding shows that there is no monolithic trust model, hence the model suitable for a project depends on various factors, including the task objectives, team composition, and other related factors.
The rest of the paper is structured as follows: Section II provides a detailed background on the evolution of trust models in HRI. Section III describes our proposed trust methodology, the simulation environment, and the case study setup of our experimental design. Section IV presents the results of our comparative analysis, followed by a discussion of the findings and their implications in Section V. Finally, Section VI concludes the paper and outlines directions for future research.
@@ II-A Evolution of Trust Models in Human-Robot Collaboration
Over the past decade, several researchers have explored different approaches for introducing computational approximates of trust into HRC systems, interfaces, and algorithms. The overall goal of this research area is to computationally capture and model the different aspects of human attitudes, behaviors, and heuristics related to trust in ways that can be used to improve performance and safety in HRC. The dimensions of prior work in this area have explored a mixture of techniques for capturing, quantifying, or predicting different trust-approximate variables.
Guo et al. [27] highlighted that trust-enabled human-robot teaming often relies on using trust models to operationalize and transform different ideals of trust into computational approaches that support human-robot teaming. They introduced the reverse psychology trust-behavior model and compared its performance against prior existing models to investigate its performance. Their work further proposed a trust-seeking reward function that balances team performance and trust preferences. Chen et al. [28] proposed trust-POMDP as a means of introducing trust variables into human-robot interaction in ways that improve performance. They tested the performance of their approach using simulation and real-life experiments and found that their approach of factoring in trust improves human-robot team performance. Yemini et al. [29] introduced a mathematical framework for determining teaming consensus, especially where there are malicious agents. They used stochastic trust values for probabilistic measures of trustworthiness between agents in a distributed consensus system. Findings from their research revealed that trust-based approaches played a significant role in enhancing and improving the security and reliability of distributed systems.
Other researchers have also explored trust using different models, including trust-aware motion planning for HRC systems [30], examining the impact of using the Theory of Mind for modeling trust-aware robot policy for HRC [31], investigating the conditions under which robots can trust human wayfinding information for navigational guidance [32], examining approaches for eliciting trust levels during live interaction in HRC [33], and highlighting areas where trust plays a crucial role in robot proficiency by foregrounding to robots the areas where their capabilities are better matched with a given task and vice versa [34]. Relevant here is that although several prior works have explored trust in HRC, very few research have explored trust in the context of complex and sometimes heterogenous MH-MR teams.
@@ II-B Factors Influencing Trust Dynamics
Hudspeth et al. [35] conducted a study to investigate the impact of interfaces on human-robot trust collaboration in physical environments. Findings from their study revealed that interfaces and feedback provided through them significantly impact levels of trust in HRC. Luebbers et al. [36] highlighted the role of human memory recall of a robot’s previous performance on trust in HRC, highlighting that a last memory of a robot failing a task leads to increased distrust of humans on robot competence. They conducted an experiment with 53 participants to test this assertion further. Findings from their study validated concerns about the impact of recency bias on human trust in robot competence. Hald and Rehm [37] used several questionnaires, including human-robot trust questionnaires, to elicit information about ways of improving the usability of HRC systems. Findings from their research using such trust questionnaires in systems led to increased usability and high trust in system performance. Rehm et al. [38] highlighted the ways in which trust plays a role in human-robot interaction in social contexts, especially with respect to users relying on or rejecting the assistance of robot-assisted shopping. Findings from their research showed that positive rejection increases user trust in the system.
Furthermore, Wang et al. [39] investigated the value of using trust for role arbitration in situations where control is alternated between humans and robots in ways that optimize team performance. They introduced a quantitative approach for modeling trust and self-confidence in HRC. Findings from their research revealed that their approach delivered superior results in HRC in achieving a task. Xu et al. [40] also investigated the approaches for enhancing trust recognition in HRC using precise approaches. They introduced an EEG-based approach that includes an EEG Vision Transformer and EEGTrust dataset for detecting and measuring levels of human trust during HRC. Findings from their research revealed that their approach produced significant results that outscored other baseline approaches. They further made their code and dataset available to facilitate this genre of research.
III Methods
@@ III-A Trust Models
We employed different trust models for our experiment, including the no-trust baseline model, Monir et al. [23], Xu and Dudek Model [24], Guo and Yang Model [25], and the Expectation Confirmation Model (ECM), which we contribute in through this research. Overall, our objective was to examine the impact of the different trust models on task allocation in the context of MH-MR. We briefly describe each model below.
@@ III-A1 No-trust baseline model
We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts and to compare its performance to other trust-enabled models.
@@ III-A4 Guo and Yang Trust Model
III-A5 Expectation Confirmation Trust (ECT) Model
@@
The ECT model addresses a critical gap by introducing trust models for MH-MR teams and explicitly incorporating the dynamic nature of human expectations in collaborative scenarios. Unlike previous single-human robot trust models that often treat trust as a static or slowly evolving parameter, the ECT model recognizes that human trust in robotic teammates is continuously shaped by the interplay between prior expectations and observed performance. This viewpoint is crucial in complex, high-stakes environments where team compositions and task demands change rapidly. By grounding the model in Expectation Confirmation Theory [26], an established framework from consumer psychology, we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [41, 42]. This interdisciplinary approach allows us to capture nuanced aspects of trust formation and maintenance in new ways.
The ECT model includes several components as illustrated in Fig. 2, including initial trust, which is often based on the assumption of a robot’s capability and perception; performance evaluation, which is based on the actual performance of a robot on a task; trust updating mechanisms that seek to match initial expectations to performance, and team-trust aggregation function that seeks to determine the overall team trust on a robot capability as a means of measuring and determining the final trust score for task allocation. Our model outline is presented in detail in Table I.
@@
             Concept                       Definition                       Formula
@@
[…]
Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1r}$
@@
Trust Update Function Trust in robot $r$ is updated based on the performance comparison ($\Delta_{ir}$) using a non-linear function to avoid drastic trust shifts due to isolated events. $T_{ir}=T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\alpha$ is the learning rate, determining how quickly trust adjusts, $\beta$ is the scaling factor that controls the influence of $\Delta_{ir}$ on trust change, and the hyperbolic tangent function ($\tanh$) ensures changes in trust are gradual and bounded.
Trust Decay Trust diminishes over time in the absence of interactions, accounting for the fading impact of past experiences. $T_{ir}=(1-\gamma)\times T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\gamma$ is the decay rate, indicating how quickly trust diminishes without recent interactions, and $T_{ir}=\min(\max(T_{ir},0),1).$
Trust Facets Trust is multi-dimensional, incorporating aspects like competence, ability, and dependability. $T_{ir}=w_{1}T_{\text{com},ir}+w_{2}T_{\text{abi},ir}+w_{3}T_{\text{dep},ir}$
where $w_{1},w_{2}$, and $w_{3}$ are weights reflecting the relative importance of each trust component.
[…]
@@
Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust are assigned simpler tasks or monitored more closely.
Team Trust Aggregation Aggregates individual trust assessments to derive a collective trust measure for each robot. $T_{\text{team}r}(t)=\frac{1}{m}\sum_{i=1}^{m}T_{ir}(t)$
which provides an overall trust level that reflects the team’s collective confidence in each robot.
[…]

### v2

# Modeling and Evaluating Trust Dynamics in Multi-Human Multi-Robot Task Allocation

@@ Abstract
Trust is essential in human-robot collaboration, particularly in multi-human, multi-robot (MH-MR) teams, where it plays a crucial role in maintaining team cohesion in complex operational environments. Despite its importance, trust is rarely incorporated into task allocation and reallocation algorithms for MH-MR collaboration. While prior research in single-human, single-robot interactions has shown that integrating trust significantly enhances both performance outcomes and user experience, its role in MH-MR task allocation remains underexplored. In this paper, we introduce the Expectation Confirmation Trust (ECT) Model, a novel framework for modeling trust dynamics in MH-MR teams. We evaluate the ECT model against five existing trust models and a no-trust baseline to assess its impact on task allocation outcomes across different team configurations (2H-2R, 5H-5R, and 10H-10R). Our results show that the ECT model improves task success rate, reduces mean completion time, and lowers task error rates. These findings highlight the complexities of trust-based task allocation in MH-MR teams. We discuss the implications of incorporating trust into task allocation algorithms and propose future research directions for adaptive trust mechanisms that balance efficiency and performance in dynamic, multi-agent environments.
@@ I Introduction
Trust is crucial in any human–robot collaboration (HRC) system, as it underpins effective coordination, fosters shared situational awareness, and minimizes misjudgments or misuse of robotic capabilities [1]. In HRC, trust has been defined in various ways, including as the belief that a human or robot teammate will perform as expected in a given task scenario [2] and as the assurance that an agent possesses sufficient capability to complete a task in uncertain and dynamic environments [3].
A key challenge lies in effectively modeling trust and leveraging it for improved collaboration. Prior research has employed various computational strategies to model trust, including probabilistic [4], non-probabilistic [3], and graph-based approaches [5], each capturing different dynamic and behavioral aspects of trust. However, despite these efforts to model trust in dyadic HRC, limited research has focused on trust dynamics within multi-agent human–robot teams [6].
Multi-human, multi-robot (MH-MR) teams [7, 8, 9, 10] are emerging as a promising paradigm for large-scale, high-stakes scenarios such as disaster response, environmental surveillance, and military operations. These teams leverage the complementary strengths of multiple humans and robots to enhance coordination, efficiency, and adaptability. Compared to traditional HRC, MH-MR interactions are more dynamic, occurring in diverse patterns such as human–human, human–robot, and hierarchical chain structures. As a result, traditional trust models may not be effective, as they often fail to account for the complex, multi-agent dependencies, shifting team compositions, and evolving situational contexts inherent in MH-MR collaboration.
@@
This research first examines how traditional HRC trust models [11, 5, 12, 13] perform in MH-MR teams. Through this investigation, we aim to emphasize the significance of trust in MH-MR teams and offer insights into the strengths and limitations of existing trust modeling approaches in these complex, collaborative settings.
We then introduce the Expectation Confirmation Trust (ECT) Model, a novel trust framework inspired by expectation confirmation theory [14], and adapt its core principles to model trust in MH-MR environments. Specifically, as illustrated in Fig. 1, the ECT model posits that trust emerges from the interplay between prior expectations of performance and actual performance outcomes, with a particular focus on how humans perceive robot performance. Its dynamic nature, along with the ability to pool trust across different team members, enhances overall team performance while facilitating cross-learning, where lower-performing agents can learn from the best-performing ones. By examining the divergence between expected and observed outcomes, the model provides a structured approach to trust calibration in dynamic, multi-agent interactions.
The main contributions are summarized below:
  • We investigate trust dynamics in the multi-human, multi-robot (MH-MR) context, an area that remains largely unexplored. Specifically, we analyze the performance of traditional HRC trust models in MH-MR settings.
  • We introduce the ECT model, a more dynamic and efficient approach for trust modeling in MH-MR task allocation. The ECT model leverages expectation confirmation theory to better capture trust adaptation and decision-making in multi-agent teams.
  • Through scalable experiments across varying team sizes: from 2 humans and 2 robots to 10 humans and 10 robots, we demonstrate the benefits of the ECT model in improving overall team performance. Specifically, we show how ECT facilitates a better trade-off between task efficiency, resource allocation, and adaptability in dynamic MH-MR environments.
@@ II-A Evolution of Trust Models in Human-Robot Collaboration
Over the past decade, researchers have developed various computational approaches to approximate trust in various HRC scenarios, like industrial collaboration [15, 16] and assistive care [17, 18]. The overarching goal of this research is to model human attitudes, behaviors, and heuristics related to trust in ways that enhance team performance, safety, and adaptability. Prior work in this area has explored diverse techniques to capture, quantify, and predict trust-related variables.
Guo et al. [19] emphasized that trust-enabled human-robot teaming relies on computational trust models to operationalize and transform abstract trust concepts into actionable mechanisms that enhance collaboration. They introduced a reverse psychology trust-behavior model and compared its effectiveness against existing models, proposing a trust-seeking reward function to balance team performance and trust preferences. Similarly, Chen et al. [20] introduced trust-POMDP, integrating trust variables into decision-making processes to optimize human-robot interaction. Their approach, validated through simulation and real-world experiments, demonstrated improved team performance when factoring trust into HRC systems.
Other studies have explored trust modeling from different perspectives. Yemini et al. [21] developed a mathematical framework for consensus in multi-agent teams, particularly in environments with malicious agents, utilizing stochastic trust values for probabilistic trust assessment. Their findings highlighted the role of trust in enhancing security and reliability in distributed systems. Yu et al. [22] investigated trust-aware motion planning for HRC, while another study by the same authors [23] examined the application of Theory of Mind in trust-aware robot policy learning. Additionally, researchers have explored how trust influences wayfinding guidance, live trust elicitation in HRC [24], and adaptive trust-based task allocation where robots assess their capabilities against task requirements.
While these studies have significantly advanced trust modeling in HRC, research on trust within complex, heterogeneous MH-MR teams remains scarce. Given the increasing prevalence of large-scale MH-MR deployments in high-stakes environments, there is a critical need to extend trust modeling frameworks beyond single-human, single-robot interactions to account for dynamic, multi-agent collaboration and adaptive trust calibration in MH-MR contexts.
@@ II-B Factors Influencing Trust Dynamics
Hudspeth et al. [25] found that interface design and feedback significantly influence trust levels in physical HRC. Luebbers et al. [26] demonstrated that recent failures lead to increased distrust due to recency bias. Hald and Rehm [27] showed that integrating trust questionnaires improves usability and system trust, and found that positive rejection in robot-assisted shopping can increase trust. Beyond cognitive and social factors, Wang et al. [28] introduced a trust-driven role arbitration model to optimize control transitions between humans and robots. Xu et al. [29] proposed an EEG-based trust recognition framework, leveraging physiological signals for real-time trust assessment [30].
While these studies provide valuable insights into trust modeling in single-human, single-robot interactions, they do not generalize well to MH-MR collaboration, where trust must be dynamically pooled, distributed, and updated across multiple agents. Unlike prior works that focus on individual trust quantification or role arbitration, our proposed ECT Model is inspired by expectation confirmation theory, which explicitly accounts for the discrepancy between expected and observed performance across multiple human and robot agents. Unlike static models that assume trust is an individual attribute, our model dynamically adjusts trust at the team level, enabling trust redistribution and cross-learning between high- and low-performing agents.
III Methodology
@@ III-A Trust Models
We consider different trust models including the no-trust baseline model, Monir et al. [11], the Xu and Dudek Model [5], the Guo and Yang Model [12]), the Ali et al. Model [13], the Guo et al. Model [6], and our Expectation Confirmation Trust model (ECT), which we contribute through this research. Overall, our objective was to examine the impact of the different trust models on task allocation in the context of MH-MR. We briefly describe each model below.
@@ III-A1 No-trust baseline model
We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts.
@@ III-A4 Guo and Yang Trust Model
III-A5 Ali et al. Trust Model
Ali et al. introduce an artificial trust measure for an agent $a$ performing a task $\gamma$ at time $t$ by comparing the task’s requirements $\bar{\alpha}_{i}$ to the agent’s current belief bounds $\ell_{i}^{a}(t)$ and $u_{i}^{a}(t)$ for each capability dimension $i=1,\dots,n$. The resulting “one-shot” trust value $\tau_{a,\gamma}(t)$ is given by
$\tau_{a,\gamma}(t)\;=\;\prod_{i=1}^{n}\psi\bigl{(}\bar{\alpha}_{i}\bigr{)},$ (8)
where the function $\psi(\bar{\alpha}_{i})$ compares $\bar{\alpha}_{i}$ against the current lower and upper bounds:
[…]
Hence, if the task requirement $\bar{\alpha}_{i}$ falls below the agent’s believed lower bound for capability $i$, trust on that dimension is $1$; if it exceeds the agent’s believed upper bound, trust on that dimension is 00; and otherwise it scales linearly in between. After each interaction (success/failure), these bounds $\ell_{i}^{a}(t)$ and $u_{i}^{a}(t)$ can be updated accordingly (e.g., expanded upon successful performances or shrunk upon failure).
III-A6 Guo et al. Trust Model
Guo et al. define trust $t_{a,b,k}$ as a Beta-distributed random variable that evolves over discrete interaction steps $k$. For agent $a$ to trust agent (or robot) $b$, they maintain Beta parameters $(\alpha_{a,b,k},\beta_{a,b,k})$ and update them after each direct or indirect experience:
$t_{a,b,k}\;\sim\;\text{Beta}\Bigl{(}\alpha_{a,b,k},\,\beta_{a,b,k}\Bigr{)}.$ (10)
The parameters $\alpha_{a,b,k}$ and $\beta_{a,b,k}$ typically increase based on positive and negative experiences, respectively. Concretely, if $p^{k}_{b}$ indicates a success/failure term for the $k$-th interaction (e.g., $p^{k}_{b}=1$ for success, $=0$ for failure), Guo et al. add coefficients $s_{a,b}$ and $f_{a,b}$ to reflect how strongly each experience changes $\alpha$ or $\beta$:
$\alpha_{a,b,k}\;=\;\alpha_{a,b,k-1}\;+\;s_{a,b}\,\bigl{(}\!\text{positive counts}\bigr{)},$ (11)
$\beta_{a,b,k}\;=\;\beta_{a,b,k-1}\;+\;f_{a,b}\,\bigl{(}\!\text{negative counts}\bigr{)}.$ (12)
Afterward, the trust value $t_{a,b,k}$ is taken as the mean (or another statistic) of the updated Beta distribution, e.g.
$t_{a,b,k}\;=\;\frac{\alpha_{a,b,k}}{\alpha_{a,b,k}+\beta_{a,b,k}}.$
More variations may also factor in indirect experiences or testimonies from other agents (i.e., other humans/robots) by incrementing $\alpha_{a}$ or $\beta_{a}$ based on second-hand observations.
III-A7 Expectation Confirmation Trust (ECT) Model
@@
The ECT model addresses a critical gap by introducing trust models for MH-MR teams and explicitly incorporating the dynamic nature of human expectations in collaborative scenarios. Unlike previous single-human robot trust models that often treat trust as a static or slowly evolving parameter, the ECT model recognizes that human trust in robotic teammates is continuously shaped by the interplay between prior expectations and observed performance. This viewpoint is crucial in complex, high-stakes environments where team compositions and task demands change rapidly. By grounding the model in Expectation Confirmation Theory [14], we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [31, 32].
As illustrated in 3, the ECT model consists of several components: (1) initial trust, based on assumptions of a robot’s capability and perception; (2) performance evaluation, which assesses a robot’s actual task performance; and (3) trust updating mechanisms that align initial expectations with observed performance. Our model outline is presented in detail in Table I and described below.
In this work, we implemented the ECT model by incorporating multi-faceted trust, adaptive learning rates, task-specific histories, and partial/conditional decay. Our overall goals were (i) to evaluate robot performance against evolving expectations, (ii) to update trust accordingly, and (iii) to decide if a given agent should attempt or decline a task under different contexts based on their score. Below, we detail each key component.
@@
             Concept                       Definition                       Formula / Note
@@
Initial Expectation ($E_{0r}$) Initial expectations for robot $r$ are set based on the task’s complexity, the environment, and the robot’s (initial) capabilities. Our model (ECT) allows skill to improve over time (cross-agent learning). $E_{0r}=g(\mathrm{complexity},\mathrm{robot\_skill})$ Our implementation example: $E=\mathrm{skill}\times(1.2-0.1\times(\mathrm{complexity}-1))$
Performance Evaluation ($P_{ir}$) After the robot executes a task ($i$), its performance is evaluated using metrics such as completion time, accuracy, and adaptability.
Successful completion can increase the robot’s skill levels in ECT, thus influencing future $E_{(i+1)r}$.
@@
Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1,r}$
@@
Trust Update Function Trust in robot $r$ is updated based on $\Delta_{ir}$ using a non-linear (tanh) function.
Note: An adaptive learning rate $\alpha_{\mathrm{dyn}}$ that increases with consistent success and decreases otherwise. Partial or selective decay (smaller if recent performance is consistent).
$T_{ir}=T_{i-1,r}+\alpha_{\mathrm{dyn}}(\text{consistency})\,\times\,\tanh\bigl{(}\beta\times\Delta_{ir}\bigr{)}$
where $\alpha_{\mathrm{dyn}}$ depends on recent performance variance/success rates, $\beta$ controls the influence of $\Delta_{ir}$ on trust change, and $\tanh$ ensures changes are bounded.
Trust Decay Trust diminishes if no recent successes, but selective decay reduces how much it fades when performance is consistent. $T_{ir}=\bigl{(}1-\gamma_{\mathrm{eff}}\bigr{)}\times T_{i-1,r}\;+\;\alpha_{\mathrm{dyn}}\,\tanh\bigl{(}\beta\times\Delta_{ir}\bigr{)},$
where $\gamma_{\mathrm{eff}}$ is the effective decay rate modulated by recent performance;
$T_{ir}=\min\bigl{(}\max(T_{ir},0),1\bigr{)}.$
Trust Facets Trust is multi-dimensional, e.g., competence, reliability, and adaptability, each updated via the above logic.
$T_{ir}=w_{1}\,T_{\mathrm{com},ir}\;+\;w_{2}\,T_{\mathrm{rel},ir}\;+\;w_{3}\,T_{\mathrm{ada},ir},$
where $w_{1},w_{2},$ and $w_{3}$ reflect each facet’s importance.
Task Allocation Decision Updated trust guides which tasks the robot attempts. Enhanced ECT also considers task-switching cost and distance/skill factors so that a more suitable agent may handle the task $A_{i+1,r}=f\bigl{(}T_{ir},\,\text{task priority},\,\text{team workload},\\[-2.15277pt] \text{switching cost},\,\text{distance factor}\bigr{)}$
@@
Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust handle simpler tasks.
Multi-Faceted Trust
Each agent $r$ maintains trust values across three distinct facets: competence, reliability, and adaptability. In vector form, we denote:
$T_{r}(t)\;=\;\bigl{(}T_{\mathrm{comp},r}(t),~{}T_{\mathrm{rel},r}(t),~{}T_{\mathrm{ada},r}(t)\bigr{)}.$
Each facet starts at an initial default trust (e.g. $0.7$). (In real-world this would be the human trust rating scores). When an agent attempts a task with complexity $c$ and obtains a performance outcome (reward) $P_{ir}\in[0,1]$, each trust facet is updated based on (i) how the actual performance compares with prior expectations, (ii) a non-linear $\tanh(\cdot)$ update, and (iii) partial decay that depends on recent consistency.
Expectations & Task Histories
For each task type (e.g. “hazard detection”) or attribute, the model tracks (a) a “task-specific trust” vector, (b) an expected performance $E_{i-1,r}$, and (c) a reward history $\{\,P_{1r},\dots,P_{i-1,r}\}$. This history is used to measure consistency, typically quantified by the variance or standard deviation of recent outcomes. A small variance (i.e. stable, repeated successes) raises the adaptive learning rate $\alpha_{\mathrm{dyn}}$ for that facet, while inconsistent or sporadic performance reduces it.
Performance Comparison
When an agent completes a task $i$, its actual performance $P_{ir}$ (a scalar reward) is compared to the prior expectation $E_{i-1,r}$. We define the “performance difference” as
$\Delta_{ir}\;=\;P_{ir}\;-\;E_{i-1,r}.$
If $\Delta_{ir}>0$, the robot is exceeding the expected baseline; if $\Delta_{ir}\leq 0$, performance is at or below expectation.
Trust Update
After the difference $\Delta_{ir}$ is computed, we perform a tanh-based update for each facet $f\in\{\mathrm{comp},\mathrm{rel},\mathrm{ada}\}$:
$T_{f,r}(t)\;\leftarrow\;T_{f,r}(t^{-})\;+\;\alpha_{\mathrm{dyn}}\,\tanh\bigl{(}\beta\,\Delta_{ir}\bigr{)},$ (13)
where $T_{f,r}(t^{-})$ is the facet’s trust just before the update, $\beta$ scales the influence of $\Delta_{ir}$, and $\alpha_{\mathrm{dyn}}$ is the adaptive learning rate, given by
$\alpha_{\mathrm{dyn}}=\alpha_{\mathrm{base}}\times\Bigl{(}1+\kappa_{\mathrm{cons}}\times\mathbbm{1}\{\text{consistent success}\}\Bigr{)}\\ \times\bigl{(}1+\mathrm{varPenalty}\bigr{)}^{-1}.$ (14)
Here, $\alpha_{\mathrm{base}}$ is a small constant (e.g. $0.05$), $\kappa_{\mathrm{cons}}>0$ is an added boost if the last few rewards were consistently above $0.7$, and $\mathrm{varPenalty}$ is proportional to the recent outcome variance. The $\tanh(\cdot)$ bounds the trust increment to avoid runaway growth.
Selective Decay
Finally, a selective decay factor is applied:
$T_{f,r}(t)\;\leftarrow\;T_{f,r}(t)\times\bigl{(}1-\gamma_{\mathrm{eff}}\bigr{)},$ (15)
where
$\gamma_{\mathrm{eff}}\;=\;\gamma_{0}\,\times\,(1-\mathrm{consistency})\,\times\,\delta(\Delta_{ir}).$
In our experiment $\gamma_{0}$ is a baseline decay rate (e.g. 0.02), consistency is high if recent standard deviation is small, and $\delta(\Delta_{ir})\in\{0.5,1.0\}$ is lower for successful tasks ($\Delta_{ir}\geq 0$) and higher for failures. This ensures that if the agent performs well consistently, trust decays only minimally, but if failures occur or consistency is low, trust decays more.
Task-Allocation Overrides
Besides updating trust, the ECT model can override an agent’s normal Q-learning action whenever its trust or skill knowledge deems the task unsuitable. Specifically, a function
$\mathrm{should\_attempt\_task}(r,\,\mathrm{task})$ (16)
checks whether $(i)$ the agent’s competence/trust is sufficiently high, $(ii)$ no other agent is significantly better suited, and $(iii)$ the agent will not pay an excessive task-switching cost by abruptly changing task types. If these conditions fail, the agent may ignore the Q-learning policy and refuse the task, thereby deferring to a more suitable teammate.
Cross-Agent Skill Learning
The ECT model also increments the agent’s skill if it succeeds on a task, especially with higher complexity. For instance, success at a complex hazard task may provide a small skill bump, increasing future $E_{r}$ or the reliability facet. Over many interactions, the agent’s trust and skill can co-evolve, ensuring that consistent good performance not only raises trust but also improves the agent’s underlying capability parameter.
Thus, the ECT model integrates:
  1. 1.
    multi-faceted trust (competence, reliability, adaptability),
  2. 2.
    adaptive trust updates (via $\tanh(\Delta)$ plus variable learning rates),
  3. 3.
    selective decay that depends on recent consistency and success/failure,
  4. 4.
    task-override logic that can bypass Q-learning decisions when trust or skill is insufficient (or if a better agent is available),
  5. 5.
    incremental skill increases upon successful task completion.
These elements combined, allow trust to reflect both short-term performance outcomes and broader trends of reliability, so that the willingness of agents to tackle more complex tasks evolves over time in a team setting.